The following table shows the number of subscribers of a cable company for recent years.

Years 2013 2014 2015 2016 2017

Subscribers

in thousands

26 36 45 115 135

a) Find the correlation coefficient for the recorded data of subscriptions.

b) Predict the number of subscribers for the year 2018.
Answer #1

Question (a)

The regression output is as follow

Subscribers SUMMARY OUTPUT Years 2013 26 2014 36 2015 45 2016 115 Regression Statistics Multiple R 0.94121282 R Square 0.8858

The correlation coefficient is the Multiple R value from the regression statistic section in the above output

so correlation coefficient = 0.941213 rounded to 6 decimal places

Question (b)

The regression equation is

\hat{y} = 29.7 * x + (-59774.1)

Where \hat{y} is the predicted number of subsribers

and x is the year

Given x = 2018

\hat{y} = 29.7 * 2018 + (-59774.1)

\hat{y} = 59934.6 - 59774.1

\hat{y} = 160.5

So the predicted number of subscribers for the year 2018 = 160.5

= 161 rouded up to nearest integer